# The ML.ENERGY Benchmark: Toward Automated Inference Energy Measurement and Optimization

### v1

@@ Appendix E Limitations
The ML.ENERGY Benchmark is not without limitations. First, we note that the benchmark is not exhaustive and does not cover all possible tasks, models, and datasets. This is particularly true as time passes and new models and tasks are developed. We are aware of newer open-weight models and worthy tasks that were released after the latest iteration of the benchmark was finalized. However, we cannot add each model or task one by one incrementally as they are released, due to the prohibitive monetary cost of running the benchmark on representative hardware; rather, we collect new advances in a window of time and then mass-update the whole benchmark, accompanied by upgrades in hardware, software, and datasets. Second, the benchmark is not exhaustive in terms of hardware. We currently mainly support flagship NVIDIA GPUs, which arguably dominates the market especially when it comes to real-world generative AI services. Furthermore, we do not have access to all possible hardware configurations, nor do they always provide a way for us to measure energy consumption from software. Regardless, we are working to expand the benchmark to support more hardware configurations.
@@ Appendix F Broader Impacts
By allowing everyone to accurate measure, understand, and optimize the energy consumption of generative AI models, we believe the ML.ENERGY Benchmark can help reduce the energy consumption of generative AI models across our community. This can lead to better democratization of the technology we develop by lowering the cost and power provisioning barrier, and enhance environmental sustainability. The authors are not aware of any negative societal impacts of the benchmark.

### v2

@@ Appendix E Limitations
The ML.ENERGY Benchmark is not without limitations. First, we note that the benchmark is not exhaustive and does not cover all possible tasks, models, and datasets. This is particularly true as time passes and new models and tasks are developed. We are aware of newer open-weight models and worthy tasks that were released after the early 2025 iteration of the benchmark was finalized. However, we cannot add each model or task one by one incrementally as they are released, due to the prohibitive monetary cost of running the benchmark on representative hardware; rather, we collect new advances in a window of time and then mass-update the whole benchmark, accompanied by upgrades in hardware, software, and datasets. Second, the benchmark is not exhaustive in terms of hardware. We currently mainly support flagship NVIDIA GPUs, which arguably dominates the market especially when it comes to real-world generative AI services. Furthermore, we do not have access to all possible hardware configurations, nor do they always provide a way for us to measure energy consumption from software. Regardless, we are working to expand the benchmark to support more hardware configurations.
@@ Appendix F Broader Impacts
By allowing everyone to accurate measure, understand, and optimize the energy consumption of generative AI models, we believe the ML.ENERGY Benchmark can enhance the understanding of energy consumption of generative AI in the research community and the industry, and ultimately fuel works that optimize energy consumption. Furthermore, energy is essentially throughput per watt, which is one factor that determines the cost of running generative AI services at the infrastructure level. By optimizing energy consumption, we can reduce the cost of running generative AI services, which can help democratize access to generative AI.